RC Meeting and Run Plan, Tuesday, February 19

From HallCWiki
Jump to: navigation, search

Main | Online | RC | Run Plan | Official Website | Redmine | Index | Nav

RC Meeting Information

RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • Continued J/Psi running on setting #1-1 through swing and owl. We gained 0.9C on this setting last night. Around 4:30am, the HB threw an interlock because the current lead mass flow read back was not working. Steve will replace the read out controller this morning.
  • We are in Restricted Access with lockup before 3pm. Tasks in the Hall: fix current lead flow meter read back on HB [1], resume work on HMS Q2 to reduce ringing (didn't have time today), fix the ripple seen in the raster (Bill Gunning is already here and says it is fine from the scope so it might be a board problem)[2], back flush magnet power supply water flows [3], install water flow regulator on HMS Q2 [4].
  • Magnets and angles are set to begin J/Psi running on Phase 2 kinematics.

Current Run Plan

Notices in Effect for Shift Crew

  • Target Operators: When the alarm handler goes white, it is disconnected from the IOC. Log it and include carlin@jlab.org every time. Acknowledge the warning. If it is still white and not coming back, restart the alarm handler. Do not permanently silence the alarm handler.
  • Missing Reference Times during the replay: check in the 50k. Check in full replay, if able. If it happens, call RC. Guidance to solve it is here: https://logbooks.jlab.org/entry/3654584
  • HMS Q1 & Q2: See notices in counting house.
  • Rotation: See notices in counting house.
  • Target camera is inoperable at the moment.

Current Issues

Planned Accesses to the Hall

  • Continued work on HMS Q2.
  • Back flush other magnets in the Hall tomorrow during RF recovery.
  • Restricted Access during Day Shift tomorrow.
  • Re-attach wooden floor stop around 57deg (HMS side)

Planned Accelerator Configuration Changes

Special Requests